OpenStudy (anonymous):

Andrew wants to buy Stephanie balloons and roses for Valentine’s Day. Roses cost $2.50 each and balloons are $1.25 each. He has $15 to spend. Which of the following is a possible combination of roses and balloons? A. 1 rose and 11 balloons B. 3 roses and 7 balloons C. 5 roses and 2 balloons D. 6 rose and 1 balloon

OpenStudy (anonymous):

@tester97

OpenStudy (tester97):

c

OpenStudy (anonymous):

C!?

OpenStudy (anonymous):

i was going to say C :)

OpenStudy (anonymous):

Lets call the number of roses 'R' and the number of balloons 'B' From the question we know that: \[($2.50 \times R) + ($1.25 \times B) = $15 \] The easiest way is to just try each answer and see which one works (trial and error). If we try D we can see that 6 roses X $2.50 = $15. No money left for the one balloon so D is out. A has the same problem: 11 balloons X $1.25 = $13.75, not enough left for a rose so A is out. B has 3 roses x $2.50 = $5, might be okay. BUT, 11 balloons x $1.25 = $13.75; puts us over. Only answer left is C. We can use our equation to check: \[($2.50 \times R) + ($1.25 \times B) = $15 \]\[($2.50 \times 5) + ($1.25 \times 2) = $15 \]\[$12.50 + $2.50 = $15 \]\[$15= $15 \] Checks.